OpenStudy (anonymous):

x-1/x over x+1/x eliminate the complex fractions

OpenStudy (anonymous):

\[\frac{x-1}{x}\times \frac{x}{x+1}\] is a start

OpenStudy (anonymous):

i.e. divide by "multiplying by the reciprocal" just like with number then cancel the \(x\) top and bottom

OpenStudy (anonymous):

as my hero @Hero would write \[\frac{x-1}{\cancel{x}}\times \frac{\cancel{x}}{x+1}=\frac{x-1}{x+1}\]

OpenStudy (anonymous):

^hey i think you misread the question

hero (hero):

@Gabysolis49, did you check your previous question?

OpenStudy (anonymous):

x-1/x over x+1/x ..isnt it (x-1/x)/(x+1/x) ?

OpenStudy (anonymous):

Hum actually he didn't the problem I have is written differently

OpenStudy (anonymous):

okay i will help you

OpenStudy (anonymous):

hey no that's different

OpenStudy (anonymous):

(x-1/x)/(x+1/x) ..this is her problem

OpenStudy (anonymous):

you would get ( x^2-1)/(x^2+1)

OpenStudy (anonymous):

no you dont..you are not interpreting it right

OpenStudy (anonymous):

jesus you still doing this?

OpenStudy (anonymous):

wait what? authority/power > correct solution?

OpenStudy (anonymous):

i know but thats what she meant as she said

hero (hero):

@PrincetonTiger, the correct way to post it linearly if we were doing it your way would be (x - (1/x))/(x + (1/x)) To post it the way I did it: ((x - 1)/x)/((x+1)/x) @Gabysolis49 posted neither so it is open to interpretation.

OpenStudy (anonymous):

perhaps it is \[\large \frac{x-\frac{1}{x}}{x+\frac{1}{x}}\] equation editor is our friend

OpenStudy (anonymous):

Yes like that^

OpenStudy (anonymous):

@Hero please. She told you that the way you are interpreting is wrong. Humans can make mistakes and so did you. Just accept it. No point pushing your authority. It is simply rude

OpenStudy (anonymous):

kk multiply top and bottom by \(x\) and finish in one step right?

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

x-1 over x+1

OpenStudy (anonymous):

^ nope that isnt right

OpenStudy (anonymous):

i would do one step for you ..((x^2-1)/x)/((x^2+1)/x))